# Faster Model Predictive Control via Self-Supervised Initialization Learning

### v2

@@ Abstract
Optimization for robot control tasks, spanning various methodologies, includes Model Predictive Control (MPC). However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability. Prior works in speeding up the optimization have limitations on optimizing MPC running time directly and generalizing to hold out domains. To overcome this challenge, we develop a novel framework aiming at expediting optimization processes directly. In our framework, we combine offline self-supervised learning and online fine-tuning to improve the control performance and reduce optimization time. We demonstrate the success of our method on a novel and challenging Formula 1 track driving task. Comparing to single-phase training, our approach achieves a $19.4\%$ reduction in optimization time and a $6.3\%$ improvement in tracking accuracy on zero-shot tracks.
@@ I INTRODUCTION
Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc. However, MPC in real-world applications is challenging due to computational time bottlenecks [8, 9, 10, 11], particularly with complex systems featuring non-convex and non-differentiable cost functions and prolonged planning horizons. [12, 13]. Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
@@
Warm starting in robot control optimization involves initializing the solver with a closer-to-optimal solution, speeding up the iterative optimization process. The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations. As shown in Figure 2, a closer initial guess reduces the distance the optimization algorithm needs to traverse to converge to the optimal solution. By starting closer to the optimum, the solver can often avoid lengthy exploration of the solution space, leading to faster convergence and reduced computational effort. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
To provide MPC with a high quality initial guess, various approaches have been adopted. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [14, 15]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [10, 16]. However, this approach may lack zero-shot generalizability. As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase, thus causing time-consuming operations to searching for the initial guess within the buffer. Alternatively Klaučo et al. propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [17]. However, this method lacks precise warm-start guesses and struggles with complex control tasks involving long planning horizons and heterogeneous observations. Moreover, it’s limited to strictly convex quadratic programs, reducing its applicability to real-world problems with non-linear or non-differentiable objectives and dynamics.
More recent approaches have adopted supervised learning to boost the performance of model predictive optimization algorithms. [18, 19] adopts behavior cloning (BC) to learn the distribution of the initial guesses for the solver. However, supervised learning cannot directly learn to minimize the computation time of MPC. Instead, all supervised learning can do is to imitate the output of the solver. Besides, BC also suffers from covariance shift, which degrades model’s performance on zero-shot scenarios [20]. [21] used reinforcement learning (RL) to train a warm-start policy to help quadrotor track simple zig-zag trajectories. Yet, RL algorithms are too data hungry to learn an optimal policy from scratch, posing challenge to complex, long-horizon tasks [22]. Furthermore, [21] is still not directly learning to minimize the computation time of MPC. Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
@@
In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time. Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional zero-shot testing phase. The two training phase works reciprocally to each other. On the one hand, the online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimize the computation time of MPC. On the other hand, the offline BC provides online learning algorithm with a good starting point to expedite the training process. Moreover, another advantage of using a learned policy to initialize the MPC, rather than for end-to-end control, is that it preserves the original definition of the control problem. The learned policy only affects the starting point, while MPC ensures the control solution satisfies system constraints, leading to safer operation. Our key contributions are:
@@
    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods all failed.
@@
    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $19.4\%$ higher performance in optimization time and a $6.3\%$ higher performance in tracking accuracy on challenging zero-shot tracks comparing to single-phase training.
Figure 2: An illustrative demo showing how a better initial guess improves MPC optimization time.
II Preliminaries
II-A Model Predictive Control
MPC is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs over a finite horizon, H. This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior. When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function , $J$ while respecting system dynamics and constraints. Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state. The optimization process is performed again at the next state. The MPC control law is formulated as shown below in Equations (1):
[…]
subject $\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
@@ II-A Model Predictive Control
Equation (1a) formulates the objective function of the MPC. $H$ denotes the planning horizon, $t$ is the current step, and, $x_{t+i}$ is the system state at the $t+i$ step. Given the state, $x_{t+i}$ and the control input, $u_{t+i}$, the system dynamics model predicts the next state, $x_{t+i+1}$, using the dynamics model (1b). Equation (1c) and Equation (1d) are the constraints for the control inputs and states. $n_{c_{u}}$ and $n_{c_{x}}$ are the number of constraints on control inputs and states respectively.
@@ III Method
In this section, we discuss in detail the proposed two-phase training framework. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
@@
1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect $N$
@@
10:Train $\pi_{\theta}^{warm}$ with Equation (2) and $\mathcal{D}$
@@
6:     Compute accumulated $xte$ using Equation (5)
@@
8:     Calculate reward in RL using Equation (4)
@@
10:     Compute training loss $L$ in Equation (3)
@@
III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs first. The expert MPC, $\pi^{MPC}$ aims to control the agent to finish the task without relying on a warm-started initial guess for the control action sequence. At each step, $\pi^{MPC}$ observes a state $s$ and takes an all-zero vector as the initial guess and optimizes that initial guess to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). Then the state-action pair, $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in the dataset $\mathcal{D}$ (line 6). The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8). During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to ensure it achieves a high control performance by disabling the “early stop” for the expert MPC.
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [23]. Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10).
BC is a simple yet effective algorithm to learn from demonstrations. The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$. BC learns a control policy, $\pi_{\theta}$, by minimizing the Mean Squared Error (MSE) in the demonstration set, shown in Equation (2). Our algorithm utilizes pre-collected MPC data to perform BC as a warm start for the MPC to improves the MPC’s optimization time.
@@ III-B Online Fine Tuning
Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through BC. This phase combines the best aspects of RL and Dataset Aggregation (DAgger) [24] algorithm. RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow A$ , is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
@@
In our framework, DAgger and RL address the covariance shift problem caused by offline BC. Additionally, RL not only fine-tunes the warm-start solution but also has the unique advantage of directly optimizing MPC computation time. We choose to leverage Twin Delayed Deep Deterministic Policy Gradient (TD3) [25] as the RL algorithm to fine-tune our pre-trained trajectory prediction model. The trajectory prediction model acts as a policy network in the TD3 framework. DAgger is integrated as a term in actor loss in the RL training, as shown in Equation (3). $B$ is the batch size and $Q_{\theta}$ is the critic network. $L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. Note that DAgger is only added when the quality of the initial guess is lower than a certain threshold. Because we only need to query the expert MPC when the initial guess is of low quality and calling the expert MPC at each step could be time-consuming. $\lambda$ is the weight coefficient between the RL loss and imitation loss.
@@
In our driving task, the reward at each step of the RL training is shown in Equation (4). The first term is the negative MPC optimization time, and the second term is the negative of the Cross Track Error ($xte$) over the planning horizon, $H$. The $xte$ is defined in Equation (5). At each step, $\pi_{\theta}^{warm}$ will output an action sequence of length $H$ based on the current state vector, $s_{t}$. Then, the dynamics model, $M_{dynamics}$, is used to calculate the future positions of the agent. The $xte$ is computed between the reference trajectory and the future positions of the vehicle, $pos_{i}^{car}$. The reward obtained at each step directly influences $L_{RL}$ by contributing to the calculation of the advantage estimate, which measures the discrepancy between the observed reward and the expected value of the state-action pair. This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (4)
$xte=distance(pos^{car}_{t},\mathbf{p}_{i}(x_{i},y_{i})^{closest})$ (5)
This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by maximizing the tracking accuracy. This helps to address the sub-optimality problem caused by BC.
IV EXPERIMENTS
In this section, we will discuss in detail the setup of the Formula 1 track domains, followed by the MPC design in this testing domain. Additionally, we will present experimental results from both training and zero-shot tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
IV-A Experiment Setup
The experiments are done in high-speed Formula 1 tracks [26] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively. The reference trajectory is represented as a set of waypoints in the center of the track. The track size is downscaled to a 10:1 ratio to make each lap a reasonable length. The reference speed for the vehicle in the expert MPC is $10m/s$. Given the downscaled size of the track data, the vehicle could actually speed up to $100m/s$ in the 1:1 scale track. The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions.
The three tracks used for demonstration collection and training are shown in Figure 3(a). The six zero-shot tracks are not presented to the vehicle before testing to demonstrate our algorithm’s zero-shot generalizability. During training, the expert MPC is first rolled out on training tracks to collect demonstrations for offline training introduced in Section III-A. Then, the warm-start policy is fine-tuned on the same three tracks using the algorithm introduced in Section III-B.
MPC is designed to control the acceleration and the steering of the vehicle to track the reference trajectory. The dynamics model of the vehicle, $M_{dynamics}$, is shown in Equation (6). $(x_{t}^{car},y_{t}^{car})$ is the global position of the vehicle at time step $t$. $v_{t}$ and $yaw_{t}$ are the speed ($m/s$) and yaw angle ($rad$) of the vehicle at time step $t$. $a_{t}$ and $\theta_{t}^{\text{steering}}$ are acceleration ($m/s^{2}$) and steering angle ($rad$) inputs to the vehicle at time step $t$. $L$ is the wheelbase of the vehicle. Here we use Ford Mustang’s dimensions, $L=2.89m$. And $dt=0.02$ is the length of each time step which represents the duration between successive updates of control inputs and system states.
$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt$ (6a)
$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt$ (6b)
$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt$ (6c)
$v_{t+1}=v_{t}+a_{t}\cdot dt$ (6d)

### v3

@@ Abstract
Model Predictive Control (MPC) is widely used in robot control by optimizing a sequence of control outputs over a finite-horizon. Computational approaches for MPC include deterministic methods (e.g., iLQR and COBYLA), as well as sampling-based methods (e.g., MPPI and CEM). However, complex system dynamics and non-convex or non-differentiable cost terms often lead to prohibitive optimization times that limit real-world deployment. Prior efforts to accelerate MPC have limitations on: (i) reusing previous solutions fails under sharp state changes and (ii) pure imitation learning does not target compute efficiency directly and suffers from suboptimality in the training data. To address these, We propose a warm-start framework that learns a policy to generate high-quality initial guesses for MPC solver. The policy is first trained via behavior cloning from expert MPC rollouts and then fine-tuned online with reinforcement learning to directly minimize MPC optimization time. We empirically validate that our approach improves both deterministic and sampling-based MPC methods, achieving up to 21.6% faster optimization and 34.1% more tracking accuracy for deterministic MPC in Formula 1 track path-tracking domain, and improving safety by 100%, path efficiency by 12.8%, and steering smoothness by 7.2% for sampling-based MPC in obstacle-rich navigation domain. These results demonstrate that our framework not only accelerates MPC but also improves overall control performance. Furthermore, it can be applied to a broader range of control algorithms that benefit from good initial guesses.
@@ I INTRODUCTION
Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc. While gradient-based MPC is highly effective when the dynamics and cost functions are smooth and differentiable, applying MPC in real-world settings remains challenging. Computational bottlenecks often arise in systems with non-convex or non-differentiable costs and long planning horizons, where gradient-free solvers are typically employed [8, 9, 10, 11].
@@
Dynamic replanning in MPC requires fast computation to operate in real time, but standard MPC without warm-starting often fails to achieve sufficient speed in challenging domains [12, 13]. Warm starting addresses this by initializing the solver with a solution close to the optimum, reducing the search space and accelerating convergence. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
Prior works have employed heuristic strategies to warm-start MPC, but these methods often struggle with scalability and fail to adapt effectively to sudden changes in system states. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [12, 14]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store and leverage historical MPC solution [10, 15]. However, this approach fails to generalize to unseen states and scales.
Traditional machine learning approach has also been adopted to learn a warm-start for MPC, however, it suffers from inaccurate initialization. Klaučo et al. propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [16], but this only produces coarse guesses. These shortcomings highlight the need for a general warm-starting framework that directly learns to generate high-quality initial guesses, scales to complex domains, and improves solver efficiency under sharp dynamics and diverse conditions.
More recent approaches have attempted to predict effective warm-starts for model predictive optimization by either behavior cloning (BC) historical trajectories [17, 18] or training reinforcement learning (RL) policies for trajectory generation [19]. While BC is fast and straightforward to train, it cannot directly learn to minimize the optimization time of MPC. Instead, all supervised learning can do is to imitate the output of the solver. Worse yet, BC suffers from covariance shift, leading to poor generalization in zero-shot scenarios [20]. In contrast, RL is capable of directly optimizing performance objectives but requires extensive interaction data and struggles to learn from scratch [21]. As such, there remains a need to bridge these two paradigms—combining the efficiency of BC with the adaptability of RL—to effectively warm-start MPC across diverse solver families.
@@
To overcome the limitations of prior work, we propose a learning framework that trains a policy to generate high-quality initial guesses for optimization-based control algorithms (e.g., MPC). The framework combines offline behavior cloning (BC) with online reinforcement learning (RL) fine-tuning and an additional zero-shot testing phase, as depicted in Fig. 1. The two training phases work synergistically. The online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimizing the optimization time of MPC. At the same time, the offline BC provides the online learning algorithm with a good starting point to expedite the training process. Moreover, initializing MPC with a learned policy rather than replacing MPC with an end-to-end controller preserves the original control formulation. The policy only sets the starting point, while MPC enforces system constraints to ensure safer operation. We demonstrate the generality of our approach by evaluating it across both deterministic MPC solver (COBYLA [22]) and sampling-based MPC solver (MPPI [23]). Our key contributions are:
@@
    We propose a novel two-phase learning framework to warm-start optimization-based control algorithms (e.g., MPC), enhancing both efficiency and robustness.
@@
    We empirically evaluate the framework on high-speed vehicle control tasks ( deterministic MPC), achieving $21.6\%$ faster optimization and $34.1\%$ higher tracking accuracy on zero-shot Formula 1 tracks compared to single-phase training.
  3. 3.
    We further test the framework on path-planning tasks (sampling-based MPC), showing $100\%$ improved safety, $12.8\%$ higher path efficiency, and $7.2\%$ smoother steering on zero-shot tracks compared to single-phase training.
II Preliminaries - Model Predictive Control
MPC is formulated as a finite-horizon optimization problem that minimizes a cost function while satisfying system dynamics and constraints. The MPC control law is formulated in Eqs.(1). At each time step $t$, MPC observes the current state $x_{t}$ and computes an optimal sequence of control inputs $U=[u_{t},\ldots,u_{t+H-1}]$ over a planning horizon $H$ by minimizing a cost function $J$ (Eq. (1a)). The system dynamics $x_{t+i+1}=f(x_{t+i},u_{t+i})$ (Eq. (1b)) predict how control inputs influence future states. Eq. (1c) and Eq. (1d) specify the constraints for the control inputs and states. After solving the optimization, only the first control input $u_{t}$ is applied, and the process repeats at the next time step using the updated state.
$J=\underset{U}{\text{minimize}}\sum_{i=0}^{H-1}l(x_{t+i},u_{t+i})$ (1a)
$\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
@@ III Method
In this section, we discuss in detail the proposed two-phase training framework to learn a warm-start policy that reduces MPC solver runtime while maintaining tracking performance.. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
@@
1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect NN
@@
10:Train $\pi_{\theta}^{warm}$ with Eq. (2) and $\mathcal{D}$
III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs. The expert MPC, $\pi^{MPC}$, controls the agent to complete the task without a warm-started initial guess. At each step, $\pi^{MPC}$ observes a state $s$, takes an all-zero vector as the initial guess, and optimizes it to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). The state-action pair $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in $\mathcal{D}$ (line 6). The first action is then applied, transitioning the system to the next state based on environment transition $T$ (line 7–8). During data collection, $\pi^{MPC}$ runs enough iterations to ensure high control performance by disabling the “early stop.”
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [24]. Given the current vehicle state, the $\pi_{\theta}^{warm}$ predicts a sequence of actions that serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10), where BC learns a control policy $\pi_{\theta}$ from pre-collected MPC demonstrations by minimizing the mean squared error (MSE) between predicted and demonstrated actions, as shown in Eq. (2).
@@ III-A Offline Training
6:  Evaluate the quality of the planned trajectory
@@
8:  Calculate reward in RL using Eq. (8)
@@
10:  Compute training loss $L$ in Eq. (3)
@@ III-B Online Fine Tuning
Our second phase, online fine-tuning, addresses the shortcomings of BC by combining the strengths of RL and Dataset Aggregation (DAgger) [25]. RL enables us to directly optimize the MPC computation time. Meanwhile, DAgger mitigates the distribution shift problem by augmenting the training set with additional expert demonstrations.
@@
RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow A$ , is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
DAgger is integrated as a term in actor loss in the RL training, as shown in Eq. (3). $B$ is the batch size and $Q_{\theta}$ is the critic network. $L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. DAgger loss is only added when the quality of the initial guess is lower than a certain threshold. To reduce the overhead of repeatedly querying the expert MPC, $L_{imitation}$ is only applied when the initial guess falls below a certain threshold. This accelerates RL training while maintaining regularization from expert demonstrations. $\lambda$ is the weight coefficient between the RL loss and imitation loss.
@@
IV EXPERIMENTS SETUP
In this section, we discuss in detail the experimental setup for both deterministic and sampling-based MPC. For deterministic MPC, we evaluate the path-tracking problem on high-speed Formula 1 tracks [26], where minimizing optimization time is critical for maintaining performance. For sampling-based MPC, we choose the navigation problem in environments without a pre-defined reference path, requiring the vehicle to generate its own collision-free trajectory.
IV-A Deterministic MPC
IV-A1 MPC Details
The Formula 1 domain consists of three training and seven zero-shot testing tracks, as shown in Fig. 2(a) and Fig. 2(b). The training tracks are used for demonstration collection and policy learning, while the zero-shot tracks remain unseen during training to test generalization. Each track’s reference trajectory is defined by waypoints along the centerline and scaled down 10:1 to make each lap a reasonable length. The friction between the tire and the road is not considered in our dynamics model.
(a) Training tracks
(b) Zero-shot tracks
Figure 2: Training and Testing maps used in the Formula 1 track path-tracking domain.
$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt+\epsilon_{1}$ (4a)
$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt+\epsilon_{2}$ (4b)
$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt+\epsilon_{3}$ (4c)
$v_{t+1}=\begin{cases}v_{t}+a_{t}\,\Delta t+\epsilon_{4},&\text{if }v_{t}+a_{t}\,\Delta t<10,\\ 10,&\text{otherwise.}\end{cases}$ (4d)
$\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I),\quad\boldsymbol{\epsilon}=[\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}]^{\top}$ (4e)
MPC controls vehicle acceleration and steering to follow the reference trajectory using the dynamics model $M_{dynamics}$ in Eq. (4). $(x_{t}^{car},y_{t}^{car})$ denote the vehicle’s global position; $v_{t}$ and $yaw_{t}$ are its speed ($m/s$) and yaw angle (radrad); $a_{t}$ and $\theta_{t}^{\text{steering}}$ represent acceleration ($m/s^{2}$) and steering angle (radrad). The wheelbase is L=2.89mL=2.89~m, and the time step is dt=0.02sdt=0.02~s. Gaussian noise $\epsilon_{1}$ - $\epsilon_{4}$ with $\sigma=0.01$ is added to model uncertainty.